OpenStudy (anonymous):

@GoldPhenoix @callie2240 -3x+y=30 -x-5y=10

OpenStudy (anonymous):

-3x+y=30 -3x y=-3x+30 Now -x-5y=10 -x -5y=-x+10 /-5 y=-x/-5-2 Now graph it go here http://www.wolframalpha.com/input/?i=plot+y%3d-3x%2b30+y%3d-x%2f-5-2&incTime=true thats your graph :D

OpenStudy (anonymous):

i told u i don't need to graph it.. i just need the work

OpenStudy (anonymous):

lol well you got the answers

OpenStudy (anonymous):

do you need the points ??

OpenStudy (anonymous):

i just need x and y.. thats all

OpenStudy (anonymous):

Oh then ignor the rest lol ill solve it

OpenStudy (anonymous):

-3x+y=30 +3x y=3x+30 Now substitue for y -3x+(3x+30)=30 30=30 There is no x so therefore there is no solution

OpenStudy (anonymous):

Understand? :D